OpenStudy (anonymous):

(4a − 0.6)^2 polynomial

jhonyy9 (jhonyy9):

so do you kmow the formula (a-b)^2 = ? than replace 4a and 0,6 and will get the right result

OpenStudy (anonymous):

1.4?

jhonyy9 (jhonyy9):

how you have got it ?

OpenStudy (anonymous):

Do what?

jhonyy9 (jhonyy9):

so do you know the formula for (a-b)^2 = ?

OpenStudy (anonymous):

Yea that is how i got 1.4

jhonyy9 (jhonyy9):

dont like writing here step by step ,please ?

jhonyy9 (jhonyy9):

(a-b)^2 =a^2 -2ab +b^2 ok ? so than now just you need replace a= 4a and b=0,6

OpenStudy (anonymous):

Ok thank you

jhonyy9 (jhonyy9):

so can you do it ?

OpenStudy (anonymous):

Yea give me a sec